The comparative balance sheets for 2018 and 2017 and the statement of income for 2018 are given below for Dux Company Additional information from Duxs accounting records is provided also. DUX COMPANY Comparative Balance Sheets Decenber 31, 2018 and 2017 s in 000s) Assets Cash Accounts receivable 2018 2017 $ 33 20 (4)So eBook 48 Less: Allovance for uncollectible accounts Dividends receivable Inventory Long-term investment Land Buildings and equipment 55 50 10 10 250 225 250 Print (25) (50 $420 369 References Liabilities $ 13 20 Accounts payable Salaries payable Interest payable Ineome tax payable Notes payable Donda payable 70 95 (21 (3) Leess Discount on bonds Sharehoiders Equity Common stook Paid-in capital-excess of par Retained earning 210 200 20 24 4547 Lessr Treasury stock (at coat) 420 $369
DUX COMPANY Income Statement For the Year Ended December 31, 2018 ($ in 000s Revenues Sales revenue s 200 Dividend revenue 3$203 Expenses Cost of goods sold Salaries expense Depreciation expense Bad debt expense Interest expense Loss on sale of building 120 25 Income tax expense Net ineone 16 178 25 Additional information from the accounting records: a. A building that originalily cost $40,000, and which was three-fourths depreciated, was sold for $7.000 b. The common stock of Byrd Corporation was purchased for $5,000 as a long-term investment. c. Property was acquired by issuing a 13%, seven-year, $30,000 note payable to the sele d. New equipment was purchased for $15,000 cash e. On January 1, 2018, bonds were sold at their $25,000 face value. t. On January 19, Dux issued a 5% stock dividend (1,000 shares). The market price of the $10 par value common stock was $14 per share at that time. g. Cash dividends of $13,000 were paid to shareholders. h. On November 12, 500 shares of common stock were repurchased as treasury stock at a cost of $8,000. Required: Prepare the statement of cash flows of Dux Company for the year ended December 31, 2018. Present cesh flows from operating activities by the direct method. (Do not round your intermediate calculations. Enter your answers in thousands. (e 10,000 should be entered as 10). Amounts to be deducted should be indicated with a minus sign.)
Answer #1

Answer :

Dux Company

Statement of Cash flow (direct method)

For the year ended December 31, 2018

$ in 000's $ in 000's
I. Cash flow from Operating activities
Cash Inflow

Cash received from customers

$ 202

Cash from Dividend

3 $ 205
Cash Outflows :

Cash paid to suppliers

132

Cash paid to employees

28

Cash paid for interest expense

6

Cash paid for income tax expense

17 (183)
Net cash from operating activities $ 22
II. Cash flow from Investing Activities
Land purchased (30)
New equipment purchased (15)
Sale of asset 7
Long term investments (5)
Net cash used by Investing activities $ (43)
III. Cash Flow from financing activities
Repurchase of treasury stock (8)
Dividends paid (13)
Bonds payable 25
Notes payable 30
Net cash flow from financing activities $ 34
Net increase in cash $ 13
Add : opening cash balance 20
Cash balance as on Dec 31,2018 $ 33

Working Notes :

1). Cash received from customers = Sales + Opening Accounts receivable - Closing Accounts receivable

= 200 + 50 - 48 = $ 202

2). Cash paid to suppliers = Cost of goods sold + Opening Accounts payable - Closing accounts payable + Ending inventory - Opening inventory

= 120 +20 - 13 + 55 - 50 = $ 132

3). Cash paid to employees = Salary expense + opening salary payable - Closing salary payable

= 25 + 5 - 2= $ 28

4). Cash paid for interest expense = 8 + 2 - 4 = $ 6

5). Cash paid for income tax = 16 + 8 - 7 = $ 17

6). Land purchased = 70 - 40 = 30